# MedReasoner: Reinforcement Learning Drives Reasoning Grounding from Clinical Thought to Pixel-Level Precision

### v2

@@ 1 Introduction
Medical imaging plays a central role in modern healthcare, where clinicians routinely examine regions of interest (ROIs) within these images to assess the health of organs and tissues (cheng2023sammed2d; lin2024stable; yan2025pgp). Consequently, precise object detection and image segmentation (often called grounding) are essential for tasks such as disease diagnosis and treatment planning (chen2021transunet; ma2024segment). To further enhance diagnostic efficiency and interpretability, medical Multimodal Large Language Models (MLLMs) have recently emerged (li2023llava; chen2024huatuogpt; xu2025lingshu). These models integrate visual perception with language interaction, allowing them to accept free-form language queries, generate high-quality responses, and even identify queried ROIs.
Despite these significant advances, a crucial limitation persists: MLLM outputs remain at the image level. To translate reasoning into visual outputs, every language reference must be grounded to a spatial location. However, while expert models (cheng2023sammed2d; yue2024surgicalsam) achieve high grounding accuracy, they rely on precise spatial prompts (e.g., bounding boxes and points). Such detailed annotations are rarely provided by clinicians in real workflows (see Fig. 1 for an example query). Recent MLLMs attempt to move beyond handcrafted prompts by coupling rich visual components (da2024segment; huang2025medseg). However, existing medical grounding pipelines are still trained in a fully supervised manner on explicitly phrased referring expressions (e.g., “segment the left lung”) (liu2023clip; koleilat2024medclip). Collecting such finely annotated data is costly and, more importantly, misaligned with real clinical queries, which are often implicit (e.g., “What can be inferred from the irregular shadow?”). Although some models can name anatomical structures, they often fail to ground them (see Fig. 1). Therefore, we need models with reasoning that can turn implicit clinical phrases into explicit spatial targets for grounding in clinical scenarios.
Existing medical visual–question answering (VQA) datasets (lau2018dataset; he2020pathvqa; liu2021slake) evaluate semantic understanding with image-level question–answer pairs but lack spatial labels. Conversely, large-scale segmentation datasets (ye2023sa; zhao2024biomedparse; li2024abdomenatlas) provide pixel-accurate masks yet omit language annotations. Neither class of dataset addresses the implicit queries that arise in real clinical practice. We have no principled way to measure whether a framework can translate implicit clinical queries into precise spatial grounding. Here, we are particularly interested in two research questions that must be addressed before implicit clinical queries can be grounded reliably:
@@
In response to RQ1, we propose U-MRG-14K, a rigorously curated dataset of 14K high-quality samples tailored to the UMRG task. U-MRG-14K is constructed from three open-source datasets. To generate semantically rich and clinically meaningful supervision, we employ GPT-4o (gpt4o) as a simulator of clinician behavior. And we design a three-stage prompting pipeline that yields high-quality QA pairs, including implicit queries, chain-of-thought (CoT) reasoning traces, and final grounding steps for each target region. Further construction details appear in Section 3.
In response to RQ2, we present MedReasoner, a reinforcement learning (RL) framework for medical reasoning and grounding. MedReasoner is decoupled into two plug-and-play components: Clinical Reasoning Module (CRM), any MLLM that reasons over implicit queries and generates lightweight spatial prompts (a bounding box plus two key points); Anatomical Segmentation Module (ASM), any model that accepts these prompts and returns a pixel-level mask. Because CRM and ASM exchange minimal geometric cues, they can be upgraded without retraining the other. Most existing grounding pipelines rely on supervised fine-tuning (SFT) with special tokens (lai2024lisa; tong2025medisee). This approach suffers from: (1) annotation hunger, it requires large, heavily annotated datasets and CoT traces are especially costly; and (2) phrase overfitting, it encourages models to echo explicit referring phrases and fails to develop genuine reasoning ability. MedReasoner solve these weaknesses through a rule-based RL training scheme that optimizes only the CRM. In each step, the CRM produces a think trace and an answer containing spatial prompts, and the frozen ASM renders a mask. Rewards for output format and spatial accuracy drive exploration, gradually aligning reasoning with precise grounding and achieving state-of-the-art performance on U-MRG-14K. As shown in Fig. 1, the RL-driven MedReasoner yields sharper grounding and more coherent reasoning than an instruction-tuned baseline, demonstrating its superiority on implicit-query grounding.
@@ 2.1 MLLMs for Medical Image Analysis
Recent advancements in MLLMs have significantly enhanced their capabilities for medical image analysis, with contributions from visual-language alignment techniques (zhu2025internvl3; wang2024healthg; Qwen2.5-VL; guo2025deepseek) . These progressions have been further extended to various medical applications, including the integration of visual expert modules into pre-trained language models (li2023llava; sellergren2025medgemma), and the unification of medical understanding and generation through heterogeneous knowledge adaptation and general foundation models (chen2024huatuogpt; lin2025healthgpt; xu2025lingshu). However, significant gaps remain in their handling of clinical complexities and crucial clinical grounding tasks, which have seen limited exploration.
@@ 2.2 Visual Grounding with Medical Reasoning
Recent MLLMs have demonstrated powerful reasoning capabilities (openai2024o1; guo2025deepseek; liu2025seg; zhu2025internvl3; Qwen2.5-VL). For visual grounding in general-purpose images, these models often leverage segmentation tools like SAM (kirillov2023segment), with methods ranging from training new tokens (lai2024lisa; ren2024pixellm) to prompting for geometric outputs (chen2024sam4mllm; uesato2022solving). However, direct application in medical scenarios is challenging due to opaque reasoning and noisy data. While some specialized works have attempted to address this (huang2025medseg; trinh2025prs; luo2024vividmed; li2025enhancing), they often struggle with the natural language found in clinical practice. Inspired by Seg-Zero (uesato2022solving), we employ reinforcement learning to generate an explicit CoT (wei2022chain). This approach enhances medical visual grounding performance while offering a transparent reasoning process, thereby increasing trust in clinical applications.
@@ 3.1 Data Generation
Most existing medical imaging datasets treat visual-grounding and VQA as separate tasks. As a result, some models support natural-language interaction without pixel-level analysis, whereas the accuracy of mainstream segmentation models hinges on the precision of supplied visual prompts. MoCoVQA (huang2025towards) attempts to unify the two tasks, yet its questions use explicit phrasing that fails to reflect the ambiguity common in routine clinical practice.
To address this gap, we construct U-MRG-14K, a medical grounding dataset centered on implicit referential expressions. U-MRG-14K is generated with GPT-4o (gpt4o) through carefully designed prompts. As shown in Fig. 2, its generation process has three stages.
@@
We collect 14K image–mask pairs from SA-Med2D-20M (ye2023sa), BiomedParse (zhao2024biomedparse), and IMIS-Bench (cheng2025interactive). We then standardize and complete the super-category labels (coarse anatomical regions) and category labels (specific organs or lesions) from the source datasets, producing a consistent and reliable taxonomy. The dataset comprises 15 super-categories and 108 categories. Table 1 provides a systematic comparison showing the advantages of U-MRG-14K over existing datasets.
@@ 4.2 Model Architecture
Enabling native pixel-level segmentation in an MLLM usually requires custom [MASK] tokens, multi-head decoders, and large collections of manual mask annotations (pi2024perceptiongpt; lai2024lisa; tong2025medisee). However, MedSAM2 (ma2025medsam2) already yields modality-agnostic masks out of the box. As shown in Fig. 3, MedReasoner decouples language reasoning from pixel-level grounding, thereby (1) mitigating phrase over-fitting to enable authentic reasoning, and (2) treating MedSAM-family models as plug-and-play components controllable by language.
@@
We employ Lingshu (xu2025lingshu) as our Clinical Reasoning Module (CRM) $\mathbf{F}_{reason}$. Given $(\mathcal{I},\mathcal{Q})$, CRM outputs a structured tuple $\texttt{<think>}\cdots\texttt{<think>}\texttt{<answer>}\mathcal{B},\mathcal{P}_{1},\mathcal{P}_{2}\texttt{<answer>}$. A bounding box $\mathcal{B}$ is often inadequate in medical images: boxes may enclose multiple organs or lesions, and their corners lack semantics for SAM-style prompters. We therefore add two key points $\mathcal{P}_{1},\mathcal{P}_{2}$ on visually distinctive regions. These enrich spatial cues at low annotation cost. To learn reliable cues without compromising linguistic competence, we train $\mathbf{F}_{\text{reason}}$ with Group Relative Policy Optimization (GRPO) (shao2024deepseekmath), using: (1) format rewards enforcing the output schema, and (2) accuracy rewards measuring spatial correctness.
@@
We instantiate the Anatomical Segmentation Module (ASM) with a frozen MedSAM2 (ma2025medsam2), denoted as $\mathbf{F}_{\text{seg}}$. The tuple $(\mathcal{B},\mathcal{P}_{1},\mathcal{P}_{2})$ produced by the CRM is fed to $\mathbf{F}_{\text{seg}}$, which transforms these coarse prompts into a high-resolution mask $\mathcal{M}$ without any task-specific fine-tuning. Freezing $\mathbf{F}_{\text{seg}}$ preserves MedSAM2’s strong zero-shot delineation ability, while allowing $\mathbf{F}_{\text{reason}}$ to concentrate on language understanding and spatial reasoning.
@@ 5.1 Experimental Settings
We conduct a comprehensive comparison across a wide range of models. For general MLLMs, we utilized GPT-4o (gpt4o), Gemini-2.5-flash (gemini2.5_flash), Qwen2.5VL-7B/72B (Qwen2.5-VL) and InternVL3-8B/78B (zhu2025internvl3). For medical-specific MLLMs, we selected MedR1-2B (lai2025med), MiniInternVL-4B (gao2024mini), MedGamma-4B (sellergren2025medgemma), HuatuoGPT-7B-Qwen2.5VL (chen2024huatuogpt), Lingshu-7B (xu2025lingshu), and Chiron-o1-8B (sun2025enhancing). For segmentation models, we chose MedSAM (ma2024segment), SAM-Med2D (cheng2023sammed2d) and MedSAM2 (ma2025medsam2). For grounding-specific models, we included SAM4MLLM (chen2024sam4mllm), VLMR1-REC-3B (shen2025vlm) and SegZero-7B (liu2025seg).
@@ 5.3 Ablation Study
This ablation study investigated how reward design influences RL training for CRM. Our SFT baseline had a low IoU of 9.15 and 2 query refusals (as Table 3 shows). RL fine-tuning drastically improved performance, eliminating all refusals. We evaluated three reward variants: Base (a hard-threshold scheme (liu2025seg)), Hard (our full reward), and Soft (IoU and pDice only). While Base removed refusals, its IoU of 15.85 was considerably lower. Our Hard reward significantly outperformed Base, increasing IoU by 15.84 points. The Soft reward variant achieved the best overall IoU of 32.42, surpassing Hard by 0.73 points, suggesting that less strict alignment fosters better exploration and more accurate grounding solutions.
@@ A.1 Referring Image Segmentation
Referring Image Segmentation (RIS) aims to segment a specific object within an image according to a natural language expression. First introduced in (hu2016segmentation), RIS addresses the need for fine-grained, language-driven visual understanding, enabling users to interact with visual systems in a more intuitive and flexible manner. Unlike traditional segmentation tasks, RIS supports flexible object descriptions, enabling intuitive interaction.
@@
This line of work narrows the language-vision gap by generating reasoning chains from text. LISA (lai2024lisa) introduces reasoning supervision via a frozen LLM and achieves strong zero-shot performance, but its token-level reasoning and reliance on supervision limit generalization. Seg-Zero (liu2025seg) improves on this with a decoupled architecture trained by reinforcement learning, enabling explicit reasoning and superior performance without annotated reasoning data.
@@
These methods improve pixel-text alignment via fine-grained feature interactions. CRIS (wang2022cris) employs contrastive learning for region-expression alignment but struggles with hard negatives and compositional queries. LAVT (yang2022lavt) enhances cross-modal encoding via early fusion, though its rigid architecture limits flexibility. GLaMM (rasheed2024glamm) enables multi-turn pixel-level grounding but requires extensive pretraining and infrastructure, hindering adaptability.
@@
This category integrates general-purpose segmentation models for enhanced control. SAM4MLLM (chen2024sam4mllm) integrates SAM with a MLLM through refinement modules, improving mask accuracy though still limited by semantic misalignment.
@@ A.2 Language-Guided Medical Image Segmentation
To support interactive and context-sensitive segmentation, some methods generate language prompts dynamically and enforce geometric consistency. FLanS (da2024segment) uses retrieval-augmented prompts with a geometry-aware model for orientation-consistent segmentations. Besides, CLIP-based models (liu2023clip) inject general language priors to support zero-shot anatomical segmentation.
@@
Another line of research enhances LGMIS by combining pretrained vision-language and segmentation models under weak or zero-shot supervision. MedCLIP-SAM (koleilat2024medclip) fuses BiomedCLIP and SAM with ScoreCAM-based supervision to achieve weakly supervised yet accurate segmentation.

### v3

@@ 1 Introduction
Medical imaging plays a central role in modern healthcare, where clinicians routinely examine regions of interest (ROIs) within these images to assess the health of organs and tissues (Cheng et al. 2023; Lin et al. 2024; Yan et al. 2025). Consequently, precise object detection and image segmentation (often called grounding) are essential for tasks such as disease diagnosis and treatment planning (Chen et al. 2021; Ma et al. 2024). To further enhance diagnostic efficiency and interpretability, medical Multimodal Large Language Models (MLLMs) have recently emerged (Li et al. 2023a; Chen et al. 2024a; Xu et al. 2025). These models integrate visual perception with language interaction, allowing them to accept free-form language queries, generate high-quality responses, and even identify queried ROIs.
Despite these significant advances, a crucial limitation persists: MLLM outputs remain at the image level. To translate reasoning into visual outputs, every language reference must be grounded to a spatial location. However, while expert models (Cheng et al. 2023; Yue et al. 2024) achieve high grounding accuracy, they rely on precise spatial prompts (e.g., bounding boxes and points). Such detailed annotations are rarely provided by clinicians in real workflows (see Fig. 1 for an example query). Recent MLLMs attempt to move beyond handcrafted prompts by coupling rich visual components (Da et al. 2024; Huang et al. 2025b). However, existing medical grounding pipelines are still trained in a fully supervised manner on explicitly phrased referring expressions (e.g., “segment the left lung”) (Liu et al. 2023; Koleilat et al. 2024). Collecting such finely annotated data is costly and, more importantly, misaligned with real clinical queries, which are often implicit (e.g., “What can be inferred from the irregular shadow?”). Although some models can name anatomical structures, they often fail to ground them (see Fig. 1). Therefore, we need models with reasoning that can turn implicit clinical phrases into explicit spatial targets for grounding in clinical scenarios.
Existing medical visual–question answering (VQA) datasets (Lau et al. 2018; He et al. 2020; Liu et al. 2021) evaluate semantic understanding with image-level question–answer pairs but lack spatial labels. Conversely, large-scale segmentation datasets (Ye et al. 2023; Zhao et al. 2024; Li et al. 2024b) provide pixel-accurate masks yet omit language annotations. Neither class of dataset addresses the implicit queries that arise in real clinical practice. We have no principled way to measure whether a framework can translate implicit clinical queries into precise spatial grounding. Here, we are particularly interested in two research questions that must be addressed before implicit clinical queries can be grounded reliably:
@@
In response to RQ1, we propose U-MRG-14K, a rigorously curated dataset of 14K high-quality samples tailored to the UMRG task. U-MRG-14K is constructed from three open-source datasets. To generate semantically rich and clinically meaningful supervision, we employ GPT-4o (OpenAI 2024a) as a simulator of clinician behavior. And we design a three-stage prompting pipeline that yields high-quality QA pairs, including implicit queries, chain-of-thought (CoT) reasoning traces, and final grounding steps for each target region. Further construction details appear in Section 3.
In response to RQ2, we present MedReasoner, a reinforcement learning (RL) framework for medical reasoning and grounding. MedReasoner is decoupled into two plug-and-play components: Clinical Reasoning Module (CRM), any MLLM that reasons over implicit queries and generates lightweight spatial prompts (a bounding box plus two key points); Anatomical Segmentation Module (ASM), any model that accepts these prompts and returns a pixel-level mask. Because CRM and ASM exchange minimal geometric cues, they can be upgraded without retraining the other. Most existing grounding pipelines rely on supervised fine-tuning (SFT) with special tokens (Lai et al. 2024; Tong et al. 2025). This approach suffers from: (1) annotation hunger, it requires large, heavily annotated datasets and CoT traces are especially costly; and (2) phrase overfitting, it encourages models to echo explicit referring phrases and fails to develop genuine reasoning ability. MedReasoner solve these weaknesses through a rule-based RL training scheme that optimizes only the CRM. In each step, the CRM produces a think trace and an answer containing spatial prompts, and the frozen ASM renders a mask. Rewards for output format and spatial accuracy drive exploration, gradually aligning reasoning with precise grounding and achieving state-of-the-art performance on U-MRG-14K. As shown in Fig. 1, the RL-driven MedReasoner yields sharper grounding and more coherent reasoning than an instruction-tuned baseline, demonstrating its superiority on implicit-query grounding.
@@ 2.1 MLLMs for Medical Image Analysis
Recent advancements in MLLMs have significantly enhanced their capabilities for medical image analysis, with contributions from visual-language alignment techniques (Zhu et al. 2025; Wang et al. 2024; Bai et al. 2025; Guo et al. 2025) . These progressions have been further extended to various medical applications, including the integration of visual expert modules into pre-trained language models (Li et al. 2023a; Sellergren et al. 2025), and the unification of medical understanding and generation through heterogeneous knowledge adaptation and general foundation models (Chen et al. 2024a; Lin et al. 2025; Xu et al. 2025). However, significant gaps remain in their handling of clinical complexities and crucial clinical grounding tasks, which have seen limited exploration.
@@ 2.2 Visual Grounding with Medical Reasoning
Recent MLLMs have demonstrated powerful reasoning capabilities (OpenAI 2024b; Guo et al. 2025; Liu et al. 2025; Zhu et al. 2025; Bai et al. 2025). For visual grounding in general-purpose images, these models often leverage segmentation tools like SAM (Kirillov et al. 2023), with methods ranging from training new tokens (Lai et al. 2024; Ren et al. 2024) to prompting for geometric outputs (Chen et al. 2024c; Uesato et al. 2022). However, direct application in medical scenarios is challenging due to opaque reasoning and noisy data. While some specialized works have attempted to address this (Huang et al. 2025b; Trinh et al. 2025; Luo et al. 2024; Li et al. 2025), they often struggle with the natural language found in clinical practice. Inspired by Seg-Zero (Uesato et al. 2022), we employ reinforcement learning to generate an explicit CoT (Wei et al. 2022). This approach enhances medical visual grounding performance while offering a transparent reasoning process, thereby increasing trust in clinical applications.
@@ 3.1 Data Generation
Most existing medical imaging datasets treat visual-grounding and VQA as separate tasks. As a result, some models support natural-language interaction without pixel-level analysis, whereas the accuracy of mainstream segmentation models hinges on the precision of supplied visual prompts. MoCoVQA (Huang et al. 2025a) attempts to unify the two tasks, yet its questions use explicit phrasing that fails to reflect the ambiguity common in routine clinical practice.
To address this gap, we construct U-MRG-14K, a medical grounding dataset centered on implicit referential expressions. U-MRG-14K is generated with GPT-4o (OpenAI 2024a) through carefully designed prompts. As shown in Fig. 2, its generation process has three stages.
@@
We collect 14K image–mask pairs from SA-Med2D-20M (Ye et al. 2023), BiomedParse (Zhao et al. 2024), and IMIS-Bench (Cheng et al. 2025). We then standardize and complete the super-category labels (coarse anatomical regions) and category labels (specific organs or lesions) from the source datasets, producing a consistent and reliable taxonomy. The dataset comprises 15 super-categories and 108 categories. Table 1 provides a systematic comparison showing the advantages of U-MRG-14K over existing datasets.
@@ 4.2 Model Architecture
Enabling native pixel-level segmentation in an MLLM usually requires custom [MASK] tokens, multi-head decoders, and large collections of manual mask annotations (Pi et al. 2024; Lai et al. 2024; Tong et al. 2025). However, MedSAM2 (Ma et al. 2025) already yields modality-agnostic masks out of the box. As shown in Fig. 3, MedReasoner decouples language reasoning from pixel-level grounding, thereby (1) mitigating phrase over-fitting to enable authentic reasoning, and (2) treating MedSAM-family models as plug-and-play components controllable by language.
@@
We employ Lingshu (Xu et al. 2025) as our Clinical Reasoning Module (CRM) $\mathbf{F}_{reason}$. Given $(\mathcal{I},\mathcal{Q})$, CRM outputs a structured tuple $\texttt{<think>}\cdots\texttt{<think>}\texttt{<answer>}\mathcal{B},\mathcal{P}_{1},\mathcal{P}_{2}\texttt{<answer>}$. A bounding box $\mathcal{B}$ is often inadequate in medical images: boxes may enclose multiple organs or lesions, and their corners lack semantics for SAM-style prompters. We therefore add two key points $\mathcal{P}_{1},\mathcal{P}_{2}$ on visually distinctive regions. These enrich spatial cues at low annotation cost. To learn reliable cues without compromising linguistic competence, we train $\mathbf{F}_{\text{reason}}$ with Group Relative Policy Optimization (GRPO) (Shao et al. 2024), using: (1) format rewards enforcing the output schema, and (2) accuracy rewards measuring spatial correctness.
@@
We instantiate the Anatomical Segmentation Module (ASM) with a frozen MedSAM2 (Ma et al. 2025), denoted as $\mathbf{F}_{\text{seg}}$. The tuple $(\mathcal{B},\mathcal{P}_{1},\mathcal{P}_{2})$ produced by the CRM is fed to $\mathbf{F}_{\text{seg}}$, which transforms these coarse prompts into a high-resolution mask $\mathcal{M}$ without any task-specific fine-tuning. Freezing $\mathbf{F}_{\text{seg}}$ preserves MedSAM2’s strong zero-shot delineation ability, while allowing $\mathbf{F}_{\text{reason}}$ to concentrate on language understanding and spatial reasoning.
@@ 5.1 Experimental Settings
We conduct a comprehensive comparison across a wide range of models. For general MLLMs, we utilized GPT-4o (OpenAI 2024a), Gemini-2.5-flash (Google 2025), Qwen2.5VL-7B/72B (Bai et al. 2025) and InternVL3-8B/78B (Zhu et al. 2025). For medical-specific MLLMs, we selected MedR1-2B (Lai et al. 2025), MiniInternVL-4B (Gao et al. 2024), MedGamma-4B (Sellergren et al. 2025), HuatuoGPT-7B-Qwen2.5VL (Chen et al. 2024a), Lingshu-7B (Xu et al. 2025), and Chiron-o1-8B (Sun et al. 2025). For segmentation models, we chose MedSAM (Ma et al. 2024), SAM-Med2D (Cheng et al. 2023) and MedSAM2 (Ma et al. 2025). For grounding-specific models, we included SAM4MLLM (Chen et al. 2024c), VLMR1-REC-3B (Shen et al. 2025) and SegZero-7B (Liu et al. 2025).
@@ 5.3 Ablation Study
This ablation study investigated how reward design influences RL training for CRM. Our SFT baseline had a low IoU of 9.15 and 2 query refusals (as Table 3 shows). RL fine-tuning drastically improved performance, eliminating all refusals. We evaluated three reward variants: Base (a hard-threshold scheme (Liu et al. 2025)), Hard (our full reward), and Soft (IoU and pDice only). While Base removed refusals, its IoU of 15.85 was considerably lower. Our Hard reward significantly outperformed Base, increasing IoU by 15.84 points. The Soft reward variant achieved the best overall IoU of 32.42, surpassing Hard by 0.73 points, suggesting that less strict alignment fosters better exploration and more accurate grounding solutions.
@@ A.1 Referring Image Segmentation
Referring Image Segmentation (RIS) aims to segment a specific object within an image according to a natural language expression. First introduced in (Hu et al. 2016), RIS addresses the need for fine-grained, language-driven visual understanding, enabling users to interact with visual systems in a more intuitive and flexible manner. Unlike traditional segmentation tasks, RIS supports flexible object descriptions, enabling intuitive interaction.
@@
This line of work narrows the language-vision gap by generating reasoning chains from text. LISA (Lai et al. 2024) introduces reasoning supervision via a frozen LLM and achieves strong zero-shot performance, but its token-level reasoning and reliance on supervision limit generalization. Seg-Zero (Liu et al. 2025) improves on this with a decoupled architecture trained by reinforcement learning, enabling explicit reasoning and superior performance without annotated reasoning data.
@@
These methods improve pixel-text alignment via fine-grained feature interactions. CRIS (Wang et al. 2022) employs contrastive learning for region-expression alignment but struggles with hard negatives and compositional queries. LAVT (Yang et al. 2022) enhances cross-modal encoding via early fusion, though its rigid architecture limits flexibility. GLaMM (Rasheed et al. 2024) enables multi-turn pixel-level grounding but requires extensive pretraining and infrastructure, hindering adaptability.
@@
This category integrates general-purpose segmentation models for enhanced control. SAM4MLLM (Chen et al. 2024c) integrates SAM with a MLLM through refinement modules, improving mask accuracy though still limited by semantic misalignment.
@@ A.2 Language-Guided Medical Image Segmentation
To support interactive and context-sensitive segmentation, some methods generate language prompts dynamically and enforce geometric consistency. FLanS (Da et al. 2024) uses retrieval-augmented prompts with a geometry-aware model for orientation-consistent segmentations. Besides, CLIP-based models (Liu et al. 2023) inject general language priors to support zero-shot anatomical segmentation.
@@
Another line of research enhances LGMIS by combining pretrained vision-language and segmentation models under weak or zero-shot supervision. MedCLIP-SAM (Koleilat et al. 2024) fuses BiomedCLIP and SAM with ScoreCAM-based supervision to achieve weakly supervised yet accurate segmentation.